# Design and Performance Simulation of the Electromagnetic Calorimeter at EicC

### v2

@@ Abstract
The electromagnetic calorimeter (ECAL) is a key detector component for precise electron and photon measurements in electron–ion collision experiments. At the Electron–Ion Collider in China (EicC), high-performance calorimetry is essential for exploring the internal structure of nucleons and studying the dynamics of quarks and gluons within quantum chromodynamics (QCD). This paper presents the optimized design and performance simulation of the EicC ECAL system. The ECAL consists of three specialized sections tailored to distinct detection environments: (1) an electron-Endcap employing high-resolution pure Cesium Iodide (pCsI) crystals, (2) a central barrel, and (3) an ion-Endcap, both adopting a cost-effective Shashlik-style sampling calorimeter with improved light yield. Each segment’s geometry and material composition have been systematically optimized through Geant4 simulations to achieve excellent energy and position resolutions as well as strong electron–pion discrimination. The simulated performance indicates that the ECAL can achieve energy resolutions of 2%/ for pCsI crystals and 5%/ for Shashlik modules, meeting the design goals of the EicC detector.
@@ 1 Introduction
The Electron-Ion Collider in China (EicC) [bib:eiccWhitePaper], proposed as a new facility at the High Intensity Heavy-ion Accelerator Facility (HIAF) [YANG2013263] in Huizhou, aims to advance our understanding of QCD. Through high-energy electron-ion collisions, it will investigate the internal structure of nucleons, the dynamics of quarks and gluons, and the fundamental nature of the strong force.
A schematic view of the EicC spectrometer is shown in the left of Fig. 1. The detector provides nearly coverage for scattered particles. Fig. 2 illustrates the momentum and pseudorapidity distributions of scattered electrons and mesons in electron–proton (ep) scattering events generated with Pythia [pythia2020], assuming 3.5 GeV electrons colliding with 20 GeV protons. The broad distributions highlight the need for uniform and high-resolution calorimetry across a wide angular range.
The ECAL of EicC is responsible for detecting electrons and photons within a pseudorapidity range of , corresponding to approximately 99% of the total solid angle. To achieve the required performance across this coverage, the ECAL is divided into three segments: the electron-endcap (e-Endcap), the central barrel, and the ion-Endcap. Each segment adopts a calorimeter technology optimized for its specific physics and geometrical constraints.
@@ 2 Electromagnetic Calorimeter Design
Based on the detection requirements across different kinematic regions at EicC, the e-Endcap, barrel, and ion-Endcap sections of the ECAL are configured to cover pseudorapidity intervals of , , and , respectively. The segments and corresponding coverages are illustrated in Fig. 1. The e-Endcap is designed for precise detection of scattered electrons, utilizing a homogeneous crystal calorimeter with excellent energy resolution. For the calorimeter in barrel and ion-Endcap, where the energy resolution is not a critical factor and the e/ identification performance will be a challenge, a Shashlik-style sampling calorimeter [bib:ecal:shashlik, Tian:2018lrh, Semenov_2020] is adopted as a compromise between the performance and the budget constraints.
@@
| Segment | Type |
| e-Endcap | CsI | -1.8 |
(-3.0, -1.0) | 4 4 | ||||||||
| Barrel | Shashlik | 0.9 | 48 (16 ) | (-1.0, 1.5) | 4 4 | ||||||||
| ion-Endcap | Shashlik | 3.5 | (1.5, 3) |
@@ 2.1 CsI Module Design
For the R&D of the e-Endcap, two types of homogeneous Cesium Iodide (CsI) crystal materials are under consideration: pure CsI (pCsI) [ATANOV2020162140] and thallium-doped CsI (CsI (Tl)) [bes3]. Both materials offer comparable costs and exhibit excellent intrinsic energy resolution, typically around 2. The main properties and differences between pCsI and CsI (Tl) are summarized in Table 2.
@@
The light yield of CsI (Tl) is approximately ten times higher than that of pCsI. However, simulation studies indicate that this increased light yield does not significantly improve the energy resolution, as the statistical uncertainty due to the photon count is already sufficiently small. One drawback of CsI (Tl) is its long scintillation decay time, which could potentially lead to event pile-up, especially under high-rate conditions. Another limitation is its relatively short attenuation length, measured to be about 30 cm for photons with a wavelength of 550 nm [bib:ecal:csi_attenuation], which can cause variations in photon collection efficiency depending on the depth of energy deposition within the crystal.
@@
The pCsI modules in the e-Endcap are designed with a projective trapezoidal geometry, featuring a total longitudinal length of 30 cm, equivalent to approximately 16 radiation lengths. Each module has a lateral dimension of 4 4 cm2 at the front face, expanding to a maximum of about 4.7 4.7 cm2 at the rear face. Scintillation photons are collected by two large-area, UV-sensitive avalanche photodiodes (APDs), model Hamamatsu S8664-1010 [bib:apd], which offer a large dynamic range and relatively high quantum efficiency in the ultraviolet region. A schematic of the pCsI module design is shown in Fig. 4.
The light yield is enhanced by coating the surface of the pCsI crystal with the wavelength-shifting (WLS) material NOL9 [bib:ecal:nol9]. NOL9 converts the ultraviolet scintillation light emitted by pCsI into yellow wavelengths, where the APDs exhibit higher quantum efficiency. The crystal module is further wrapped with Teflon tape, providing high reflectivity, particularly in the ultraviolet region. Optical silicone grease is applied between the APDs and the crystal surface to remove any air gap, as an air layer could cause total internal reflection and reduce photon collection efficiency. For mechanical stability and to maintain a light-tight enclosure, the entire assembly is housed within a carbon-fiber shell.
@@ 2.2 Shashlik Module Design
The Shashlik-style calorimeter is a sampling calorimeter consisting of alternating layers of active and absorbing materials, with wavelength-shifting (WLS) fibers penetrating through each layer to collect scintillation light. It provides moderate energy resolution and has been adopted as a cost-effective solution for the ECAL calorimeters in both the barrel and ion-endcap regions. Recent beam tests have demonstrated that the Shashlik calorimeter can achieve an energy resolution of approximately 4.5%/ [bib:ecal:nica_beam], which is sufficient to meet the requirements for high-energy photon measurement and e/ PID. The design parameters of the Shashlik calorimeter, including the sampling ratio, the number of layers, and the lateral cell size, can be adjusted to meet ECAL’s specific performance requirements. For example, the longitudinal structure primarily affects the energy resolution, whereas the lateral dimension influences the position resolution.
In the current design, the alternating layers of the Shashlik module have been carefully optimized to balance performance requirements and engineering constraints. Based on simulation studies and experience from previous Shashlik calorimeter designs [bib:ecal:shashlik, Tian:2018lrh, Semenov_2020], a configuration comprising 240 layers of 1.5 mm thick plastic scintillator as the active material and 0.35 mm thick lead plates as the absorber has been selected, the design of which has a sampling ratio of 0.33, as illustrated in Fig. 5. This configuration provides a total thickness of approximately 16 , enabling effective photon detection for energies up to 15 GeV.
@@
The Shashlik module features a lateral size of 4 4 cm2. Each module is equipped with 16 double-cladding WLS fibers, with a diameter of 1.2 mm, which penetrate longitudinally through the holes in each layer to collect both scintillation and Cherenkov photons. The collected light from the WLS fibers is read out by a 6 6 mm2 SiPM mounted at one end of the module. To keep all layers firmly compressed together, four stainless-steel rods are used to apply adequate mechanical pressure.
To enhance the light yield, a 65 m thick 3M™ ESR mirror film, with a typical reflectivity greater than 98% for visible light, is inserted between the scintillator and lead plates as a reflective layer. The ESR film is also attached to the distal end of each WLS fiber, serving as a mirror reflector. Additionally, the outer surface of the module is coated with TiO2 paint to improve light collection further and keep module light-tight.
@@
3 Simulation Result and Detector Design Optimization
To evaluate the performance of the modules, Geant4 [geant4] simulations were conducted using a 7 7 module array for both the pCsI and Shashlik calorimeters. This configuration is sufficient to contain the majority of the lateral energy deposition from electromagnetic showers. The simulations focus on both the energy deposition in sensitive materials and the photon transmission processes within the calorimeter. The evaluation and optimization of the detector design focus on several key performance metrics, including energy resolution, position resolution, and PID capability.
@@ 3.1 Energy Reconstruction and Resolution
The energy resolution of both the pCsI and Shashlik calorimeters at a fixed incident energy is obtained by fitting the energy deposition spectrum with a Gaussian function. However, for the pCsI crystal module, the energy deposition does not follow a purely Gaussian distribution due to the presence of asymmetric low-energy tails caused by energy leakage or photons escape. Instead, the distribution is well described by the Crystal Ball function [crystal_ball], which combines a Gaussian core with a power-law low-end tail. For the pCsI crystal, the energy resolution is approximated by the full width at half maximum (FWHM) divided by 2.35.
The energy resolution as a function of incident energy is fitted using a function with the electronic noise term ignored for the simulations [Fabjan2020]:
where is the stochastic term, which dominates the energy resolution at low energies, and is the constant term, typically associated with energy leakage and calibration uncertainties. The symbol denotes a quadratic sum of the two components.
The simulated energy resolution for both electrons and photons are shown in Fig. 6, illustrating three stages of evaluation: the intrinsic energy deposition (Edep), the photoelectron statistics (NPE), and the reconstructed cluster energy obtained from NPE after applying the neural-network and cellular-automata clustering algorithm [bib:ecal_algorithm]. The pCsI crystal calorimeter achieves an energy resolution of 1.76%/ for electrons, while the Shashlik calorimeter reaches 5%/. The photon energy resolution is slightly worse than that for electrons, primarily because photons have a smaller interaction cross section with matter, leading to a deeper shower start position and consequently a higher probability of energy leakage. For the Shashlik calorimeter, the energy resolution based on NPE is significantly degraded compared with the Edep case, mainly due to the non-uniformity of light collection in the scintillator, particularly its dependence on the transverse position of energy deposition.
@@ 3.2 Position Reconstruction and Resolution
The hit position in the ECAL is reconstructed using the logarithmic energy-weighted average method [bib:ecal:position], and also considers the positions of modules with energy deposition requirement. The reconstructed position is calculated as:
@@ 3.3 Electron and Pion Identification
As shown in the left panel of Fig. 8, the ratio for electrons is around value 1, while most pions have a ratio less than 1. By applying an requirement, the majority of pion contamination can be suppressed, thereby improving electron purity. However, it should be noted that a fraction of pions can undergo hadronic interactions within the calorimeter and deposit energy up to their energy. In such cases, these pions mimic the response of electrons and can contaminate the electron events, particularly when the production ratio is large.
Additional discrimination between electrons and pions can be achieved by exploiting the difference in the spatial development of their showers. In general, hadronic showers tend to be broader and exhibit a larger lateral spread compared to the more compact electromagnetic showers induced by electrons. This characteristic can be quantified using the second moment of the hit position distribution, referred to as the shower dispersion [bib:ecal:position]. The total dispersion , which accounts for both the and directions, is defined as:
@@
Both the and dispersion () requirements are applied to achieve high electron efficiency while ensuring effective pion rejection. The right panel of Fig. 8 shows the distribution of dispersion versus the ratio for electrons and pions in the Shashlik array. A circular requirement is employed to optimize hadron rejection, effectively suppressing pion contamination while maintaining high electron detection efficiency. After applying the combined and requirements, the performance reaches a pion rejection factor of about 100:1 with an electron efficiency exceeding 99%, as summarized in Table 3.
@@
In our e/ PID study, an equal number of electrons and pions with a momentum of 2 GeV/ are simulated. In realistic experimental conditions, however, the /e ratio can be much larger, potentially reaching values as high as . This highlights the stringent requirements for PID performance. A promising strategy to further enhance e/ discrimination, particularly at low momentum, is to combine the calorimeter energy-deposit information with complementary measurements from other PID detectors, such as Cherenkov and Time-of-Flight (TOF) detectors.

### v3

@@ Abstract
The Electromagnetic Calorimeter (ECAL) is a key detector component for precise measurements of electrons and photons in electron–ion collision experiments. At the Electron–Ion Collider in China (EicC), high-performance calorimetry is essential for exploring the internal structure of nucleons and studying the dynamics of quarks and gluons within quantum chromodynamics (QCD). This paper presents the optimized design and performance simulation of the EicC ECAL system. The ECAL consists of three specialized sections tailored to distinct detection environments: (1) an electron-Endcap employing high-resolution pure Cesium Iodide (pCsI) crystals, (2) a central barrel, and (3) an ion-Endcap, both adopting a cost-effective Shashlik-style sampling calorimeter with improved light yield. Each segment’s geometry and material composition have been systematically optimized through Geant4 simulations to achieve excellent energy and position resolutions as well as strong electron–pion discrimination. The simulated performance indicates that the ECAL can achieve energy resolutions of 1.76%/ for pCsI crystals and 5%/ for Shashlik modules, meeting the design goals of the EicC detector.
@@ 1 Introduction
The Electron–Ion Collider in China (EicC) [2], proposed as a new facility at the High Intensity Heavy-ion Accelerator Facility (HIAF) [20] in Huizhou, is designed to advance our understanding of quantum chromodynamics (QCD). Through high-energy electron–ion collisions, it will probe the internal structure of nucleons, with particular emphasis on the origins of nucleon spin and mass, the dynamics of quarks and gluons, and the fundamental properties of the strong interaction.
A schematic view of the EicC spectrometer is shown on the left of Fig. 1. The detector provides nearly coverage for scattered particles. Fig. 2 shows the momentum and pseudorapidity distributions of scattered electrons and mesons, while Fig. 3 presents the corresponding ratio. All figures are obtained from scattering events generated with Pythia [17], assuming 3.5 GeV electrons colliding with 20 GeV protons. The broad kinematic distributions highlight the need for uniform and high-resolution calorimetry across a wide angular range. The expected luminosity is , corresponding to an interaction rate of approximately 83.2 kHz, with an average final-state particle multiplicity of about 9 per event.
Due to the asymmetric beam energies at the EicC, the ECAL must operate effectively across a wide and region-dependent dynamic range. The ECAL is responsible for detecting electrons and photons within the pseudorapidity range , which corresponds to approximately 99% of the total solid angle. To achieve the required performance across this coverage, the ECAL is divided into three segments: the electron-endcap (e-Endcap), the central barrel, and the ion-Endcap. Each segment adopts a calorimeter technology optimized for its specific physics and geometrical constraints.
@@ 2 Electromagnetic Calorimeter Design
Based on the detection requirements across different kinematic regions at EicC, the e-Endcap, barrel, and ion-Endcap sections of the ECAL are configured to cover pseudorapidity intervals of , , and , respectively. The segments and corresponding coverages are illustrated in Fig. 1. The e-Endcap is designed for precise detection of scattered electrons, utilizing a homogeneous crystal calorimeter with excellent energy resolution. For the calorimeter in barrel and ion-Endcap, where the energy resolution is not a critical factor and the e/ identification performance will be a challenge, a Shashlik-style sampling calorimeter [4, 18, 16] is adopted as a compromise between the performance and the budget constraints.
@@
| Parameter | e-Endcap | Barrel | ion-Endcap |
| Material | pCsI | Shashlik | |
| Distance to IP [m] | 1.8 | 0.9 (radius) | 3 |
| acceptance | [-3, -1] | [-1, 1.5] | [1.5, 3] |
| Inner radius [cm] | 18 | 90 | 30 |
| Outer radius [cm] | 153 | 150 | 145 |
| Length [cm] | 30 + 5 (readout) | 49 + 11 (readout) | |
| Radiation length [] | 16 | 16 | |
| Molire radius [cm] | 3.5 | 5.02 | |
| Front size [] | 44 | 44 | |
| Rear size [] (max) | 4.674.67 | 5.86.1 | 4.64.6 |
| N layers | - | 240 | |
| Scintillator thickness [mm] | - | 1.5 | |
| Lead thickness [mm] | - | 0.35 | |
| Reflector thickness [mm] | - | 0.075 | |
| Sampling ratio | - | 0.33 | |
| N fibers (front) | - | 16 | |
| Photon detector | APD or SiPM | 66 SiPM | |
| Total modules | 3900 | 8000 | 3900 |
In addition to the geometrical requirements, the expected energy coverage of the calorimeter is determined by the underlying event energy distributions, as previously shown in Fig. 2. For the e-Endcap, the crystals are designed to measure energy deposits from 10 MeV up to 4 GeV. For the Shashlik calorimeters, the minimum detectable energy is about 25 MeV, while the maximum energy reaches about 10 GeV in the barrel region and up to 15 GeV in the ion-Endcap. These ranges define the required dynamic measurement capability of the detector and impose corresponding constraints on the detector design, including sensor response and signal readout.
@@ 2.1 CsI Module Design
For the R&D of the e-Endcap, two types of homogeneous Cesium Iodide (CsI) crystal materials are under consideration: pure CsI (pCsI) [3] and thallium-doped CsI (CsI (Tl)) [13]. Both materials offer comparable costs and exhibit excellent intrinsic energy resolution, typically around 2. The main properties and differences between pCsI and CsI (Tl) are summarized in Table 2.
@@
The light yield of CsI (Tl) is approximately ten times higher than that of pCsI. However, simulation studies indicate that this increased light yield does not significantly improve the energy resolution, as the statistical uncertainty from photon counting is already sufficiently small. One drawback of CsI (Tl) is its long scintillation decay time, which could potentially lead to event pile-up, especially under high-rate conditions. Another limitation is its relatively short attenuation length, measured to be about 30 cm for photons with a wavelength of 550 nm [11], which can cause variations in photon collection efficiency depending on the depth of energy deposition within the crystal.
@@
The pCsI modules in the e-Endcap are designed with a projective trapezoidal geometry, featuring a total longitudinal length of 30 cm, equivalent to approximately 16 radiation lengths. Each module has a lateral dimension of 4 4 cm2 at the front face, expanding to a maximum of about 4.7 4.7 cm2 at the rear face. Scintillation photons are collected by two large-area, UV-sensitive avalanche photodiodes (APDs), model Hamamatsu S8664-1010 [9], which offer a large dynamic range and relatively high quantum efficiency in the ultraviolet region. A schematic of the pCsI module design is shown in Fig. 5.
The light yield is enhanced by coating the surface of the pCsI crystal with the wavelength-shifting (WLS) material NOL9 [6]. NOL9 converts the ultraviolet scintillation light emitted by pCsI into yellow wavelengths, where the APDs exhibit higher quantum efficiency. The crystal module is further wrapped with Teflon tape, providing high reflectivity, particularly in the ultraviolet region. Optical silicone grease is applied between the APDs and the crystal surface to eliminate any air gap, as an air layer could cause total internal reflection and reduce photon collection efficiency. For mechanical stability and to maintain a light-tight enclosure, the entire assembly is housed within a carbon-fiber shell.
@@ 2.2 Shashlik Module Design
The Shashlik-style calorimeter is a sampling calorimeter comprising alternating layers of active and absorbing materials, with wavelength-shifting (WLS) fibers passing through each layer to collect scintillation light. It provides moderate energy resolution and has been adopted as a cost-effective solution for the ECAL calorimeters in both the barrel and ion-endcap regions. Recent beam tests have demonstrated that the Shashlik calorimeter can achieve an energy resolution of approximately 4.5%/ [14], which is sufficient to meet the requirements for high-energy photon measurement and e/ PID. The design parameters of the Shashlik calorimeter, including the sampling ratio, the number of layers, and the lateral cell size, can be adjusted to meet ECAL’s specific performance requirements. For example, the longitudinal structure primarily affects the energy resolution, whereas the lateral dimension influences the position resolution.
In the current design, the alternating layers of the Shashlik module have been carefully optimized to balance performance requirements and engineering constraints. Based on simulation studies and experience from previous Shashlik calorimeter designs [4, 18, 16], a configuration comprising 240 layers of 1.5 mm thick plastic scintillator as the active material and 0.35 mm thick lead plates as the absorber has been selected, the design of which has a sampling ratio of 0.33, as illustrated in Fig. 6. This configuration provides a total thickness of approximately 16 , enabling effective photon detection for energies up to 15 GeV.
@@
The Shashlik module features a lateral size of 4 4 cm2. Each module is equipped with 16 double-cladding WLS fibers, with a diameter of 1.2 mm, which penetrate longitudinally through the holes in each layer to collect both scintillation and Cherenkov photons. The light collected by the WLS fibers is read out by a 6 6 mm2 SiPM mounted at one end of the module. To keep all layers firmly compressed together, four stainless-steel rods are used to apply adequate mechanical pressure.
To enhance the light yield, a 65 m thick 3M™ ESR mirror film, with a typical reflectivity greater than 98% for visible light, is inserted between the scintillator and lead plates as a reflective layer. The ESR film is also attached to the distal end of each WLS fiber, serving as a mirror reflector. Additionally, the outer surface of the module is coated with TiO2 paint, which enhances light collection and ensures optical isolation of the module.
Simulation results indicate that a 1 GeV electron yields approximately 8000 photoelectrons at the SiPM readout. However, the actual light yield strongly depends on the light yield of the scintillator, the specific SiPM model employed, and the photon detection efficiency (PDE) at the chosen operating overvoltage. In practice, the overvoltage is often limited by the dynamic range of the readout electronics, which can lead to a reduced effective PDE.
@@
3 Simulation Results and Detector Design Optimization
To evaluate the performance of the modules, Geant4 [1] simulations were conducted using a 7 7 module array for both the pCsI and Shashlik calorimeters. This configuration is sufficient to contain the majority of the lateral energy deposition from electromagnetic showers. A uniform square particle source with the same transverse size as a single module () was used as input, allowing particles to hit the full module surface and thereby accounting for the overall impact of different incident positions on the detector performance. The simulations focus on both the energy deposition in sensitive materials and the photon transmission processes within the calorimeter. The evaluation and optimization of the detector design focus on several key performance metrics, including energy resolution, position resolution, and PID capability.
In the simulation of photon transmission processes, a detailed and realistic detector geometry is implemented in Geant4, including scintillation materials, reflective layers, wavelength-shifting materials (fibers and NOL9 painting), and optical sensors. Optical photon production and transport are fully simulated, covering scintillation light generation, reflection, absorption, wavelength shifting, and final photon detection. All material properties are taken from manufacturer datasheets or our experimental measurements, such as light yield, attenuation length, reflectivity, wavelength conversion efficiency, and photodetector quantum efficiency. Due to the large light yield and multiple internal reflections, the full optical simulation is computationally intensive and results in relatively low running efficiency.
A cluster reconstruction algorithm based on a neural-network and cellular-automata approach [7] is applied using the number of photoelectrons (NPE) information of individual modules. Adjacent modules with high NPE are grouped into a single cluster, while isolated modules with low energy deposits and no neighbours are discarded in the reconstruction procedure. This suppression of low-energy isolated hits reduces noise contributions but may introduce a small degradation in the energy resolution. This clustering algorithm is also employed in the reconstruction to identify two-photon candidates.
@@ 3.1 Energy Reconstruction and Resolution
The energy resolution of both the pCsI and Shashlik calorimeters at a fixed incident energy is obtained by fitting the energy deposition spectrum with a Gaussian function. However, for the pCsI crystal module, the energy deposition does not follow a purely Gaussian distribution due to the presence of asymmetric low-energy tails caused by energy leakage or photons escaping. Instead, the distribution is well described by the Crystal Ball function [15], which combines a Gaussian core with a power-law low-end tail. For the pCsI crystal, the energy resolution is approximated by the full width at half maximum (FWHM) divided by 2.35.
The energy resolution as a function of incident energy is fitted using a function with the electronic noise term ignored for the simulations [8]:
where is the stochastic term, which dominates the energy resolution at low energies, and is the constant term, mainly arising from shower leakage (both lateral and longitudinal) within the finite array, as well as non-uniform light collection among modules in the NPE-based results. The symbol denotes a quadratic sum of the two components.
The simulated energy resolution results for both electrons and photons are shown in Fig. 7, illustrating three stages of evaluation: the intrinsic energy deposition (Edep), the NPE statistics, and NPE statistics with applying the cluster reconstruction algorithm. The pCsI crystal calorimeter achieves an energy resolution of 1.76%/ for electrons, while the Shashlik calorimeter reaches 5%/. The photon energy resolution is slightly worse than that for electrons, primarily because photons have a smaller interaction cross section with matter, leading to a deeper shower start position and consequently a higher probability of energy leakage. For the Shashlik calorimeter, the energy resolution derived from NPE is significantly degraded compared to the Edep case, particularly in the constant term. This deterioration mainly arises from non-uniform light collection efficiency in the scintillator, especially its dependence on the transverse position of the energy deposition.
@@ 3.2 Position Reconstruction and Resolution
The hit position in the ECAL is reconstructed using a logarithmic energy-weighted centroid method [5], where the deposited energy (or NPE) in each module is used as the weight. Only modules above a certain energy threshold are included in the calculation. The reconstructed position is given by:
@@ 3.3 Electron and Pion Identification
As shown in the left panel of Fig. 9, the ratio for electrons is around the value 1, while most pions have a ratio less than 1. By applying an requirement, the majority of pion contamination can be suppressed, thereby improving electron purity. However, it should be noted that a fraction of pions can undergo hadronic interactions within the calorimeter and deposit energy up to their energy. In such cases, these pions mimic the response of electrons and can contaminate the electron events, particularly when the production ratio is large.
Additional discrimination between electrons and pions can be achieved by exploiting the difference in the spatial development of their showers. In general, hadronic showers tend to be broader and exhibit a larger lateral spread compared to the more compact electromagnetic showers induced by electrons. This characteristic can be quantified using the second moment of the hit position distribution, referred to as the shower dispersion [5]. The total dispersion , which accounts for both the and directions, is defined as:
@@
Both the and dispersion () requirements are applied to achieve high electron efficiency while ensuring effective pion rejection. The right panel of Fig. 9 shows the distribution of dispersion versus the ratio for electrons and pions in the Shashlik array. The observed structures in reflect the distinct shower distributions: corresponds to localized MIP-like pion events in a single module; corresponds to energy sharing between two adjacent modules along a single transverse axis; and () indicates the four-cell sharing pattern. A circular requirement is employed to optimize hadron rejection, effectively suppressing pion contamination while maintaining high electron detection efficiency. After applying the combined and requirements, the performance reaches a pion rejection factor of about 100:1 with an electron efficiency exceeding 99%, as summarized in Table 3.
@@
The Fig.10 shows the pion rejection and electron efficiency of the pCsI and Shashlik arrays at different momenta. To better approximate realistic experimental conditions, the ratio in the simulation is set to 10:1. It can be observed that both the electron efficiency and pion rejection generally improve with increasing energy. For the Shashlik calorimeter, however, the electron efficiency decreases slightly at higher energies. This is because high-energy pions increasingly exhibit electromagnetic shower-like behavior, making them more difficult to distinguish from electrons and degrading particle identification performance.
In realistic experimental conditions, as shown in Fig. 3, the ratio can reach magnitudes as high as , which further highlights the stringent requirements on PID performance. Several approaches can be pursued to improve the PID capability of the calorimeter system. First, further optimization of the calorimeter-based PID algorithms is necessary. Second, improvements at the detector design level, including the introduction of a preshower detector [10] and a longitudinally segmented readout scheme [12], can provide more detailed information on the early stage and longitudinal development of particle showers. Finally, effective particle identification also relies on the combined use of other PID detectors, such as Cherenkov and Time-of-Flight (TOF) systems. Such a multi-detector approach is particularly important for enhancing e/ discrimination in the low-momentum region.